# Improved Sample Complexity For Diffusion Model Training Without Empirical Risk Minimizer Access

### v5

@@ Abstract
Diffusion models have demonstrated remarkable performance in generating high-dimensional samples across domains such as vision, language, and the sciences. Although continuous-state diffusion models have been extensively studied both empirically and theoretically, discrete-state diffusion models, essential for applications involving text, sequences, and combinatorial structures, they remain significantly less understood from a theoretical standpoint. In particular, all existing analyses of discrete-state models assume access to an empirical risk minimizer. In this work, we present a principled theoretical framework analyzing diffusion models, providing a state-of-the-art sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$. Our structured decomposition of the score estimation error into statistical and optimization components offers critical insights into how diffusion models can be trained efficiently. This analysis addresses a fundamental gap in the literature and establishes the theoretical tractability and practical relevance of diffusion models.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks (Ulhaq & Akhtar, 2022; Bansal et al., 2023), text generation (Li et al., 2022), sequential data modeling (Tashiro et al., 2021), reinforcement learning and control (Zhu et al., 2023), and life sciences (Jing et al., 2022; Malusare & Aggarwal, 2024). For a more comprehensive exposition of applications, we refer readers to survey paper (Chen et al., 2024).
@@
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. Some of the key works studying the sample complexity are summarized in Table 1. A key limitation of sample complexity analyses of diffusion models done thus far is the lack of the presence of finite-time sample complexity results under reasonable assumptions. This makes the theoretical analysis of diffusion models fall short of other machine learning areas such as reinforcement Learning (Kumar et al., 2023; Gaur et al., 2024), bi-level-optimization (Grazzi et al., 2023; Gaur et al., 2025) and graphical models (Fattahi et al., 2019; Tran et al., 2019). In this work we aim to bridge that gap and obtain a sample complexity results on the same footing as results from the aforementioned areas. The iteration complexity or convergence has been studied in Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b), while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function.
@@
We note that works such as Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023); Chen et al. (2023) have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, Gupta et al. (2024) improved upon this by obtaining $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity without exponential dependence on data dimension. [^1] We note that Gupta et al. (2024) claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix A.. However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss, a significant restriction that was explicitly highlighted as an open problem in Gupta et al. (2024) itself. While this assumption is present in all prior works, it is an unrealistic assumption regardless.
@@
Zhang et al. (2024) $\tilde{O}\left(\epsilon^{-d}\right)$ Yes
Wibisono et al. (2024) $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes
Oko et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Chen et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Gupta et al. (2024)1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes
@@
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. For further details on how the sample complexity bounds are derived for Gupta et al. (2024), see Appendix A.
@@
Unconditional and Conditional Diffusion Models. Diffusion models have emerged as leading frameworks across vision, audio, and scientific domains. Foundational works such as Sohl-Dickstein et al. (2015) and Ho et al. (2020) introduced and refined Denoising Diffusion Probabilistic Models (DDPMs), enabling high-quality sample generation. Subsequent advances include improved noise schedules (Nichol & Dhariwal, 2021), score-based SDE formulations (Song et al., 2021), and efficient latent-space generation via Latent Diffusion Models (LDMs) (Rombach et al., 2022). Conditional diffusion models extend these techniques for guided generation tasks, with applications in time-series Tashiro et al. (2021), speech Huang et al. (2022), and medical imaging Dorjsembe et al. (2023). Conditioning mechanisms range from classifier-based Dhariwal & Nichol (2021) to classifier-free guidance Ho & Salimans (2022), which enabled text-to-image models like Imagen Saharia et al. (2022) and Stable Diffusion Rombach et al. (2022). Recent innovations focus on adaptive control Castillo et al. (2025), compositionality Liu et al. (2023), and multi-modal conditioning Avrahami et al. (2022).
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in Chen et al. (2022) showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b). In particular, Benton et al. (2024); Li et al. (2024b) establishes iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates Li et al. (2024a); Liang et al. (2025b); Dou et al. (2024). Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in Li & Yan (2024); Huang et al. (2024); Liang et al. (2025a). In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the errors introduced by the neural network approximation, data sampling, and optimization are also accounted for.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in Chen et al. (2023); Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023). However, these bounds are exponential in the data dimension. Further, authors of Gupta et al. (2024) use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
@@ 2 Preliminaries and Problem Formulation
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process (Øksendal, 2003), defined by the following SDE
@@
This is typically achieved using stochastic time-reversal theory (Anderson, 1982), which yields a corresponding reverse-time SDE as follows.
@@
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions. In practice, the backward process is run up to a fixed time point $t_{0}$ known as the early stopping time and not $t=0$. This is done in order to improve performance and training speed (Lyu et al., 2022; Favero et al., 2025).
The continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework (Ho et al., 2020), where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix B. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
@@ 3 Sample Complexity of Diffusion Models
Some works that analyze the convergence of score-based diffusion models, such as Chen et al. (2022); Oko et al. (2023), assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to be sub-Gaussian, making our results applicable to a significantly broader class of distributions.
@@
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses (Liu et al., 2022). Prior works such as Gupta et al. (2024) and Block et al. (2020) implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in Gupta et al. (2024) and the definition of $\hat{f}$ in Theorem 13 of Block et al. (2020)). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.
@@
This error is independent of the sampling algorithm, and describes the error due to neural network parametrization. In learning theory, it is common to treat the approximation error of a model class as a constant so that analyzes can focus on the estimation/ optimization terms dependent on the sample. This convention appears in standard excess-risk decompositions for fixed hypothesis classes (Shalev-Shwartz & Ben-David, 2014). In PAC-Bayesian analyses, approximation errors are denoted by a constant once the class is fixed (Mai, 2025). In (NTK/RKHS) analyses of neural networks, where it is assumed the target function lies in, or is well approximated by the specified function class, the misspecification error is represented as a constant term (Bing et al., 2025). In reinforcement learning algorithm analysis such as policy gradient, a task-dependent “inherent Bellman” or function-approximation error that remains constant while deriving performance rates (Mondal & Aggarwal, 2024; Fu et al., 2021; Gaur et al., 2024; Ganesh et al., 2025). Note that Gupta et al. (2024) also makes the same assumption implicitly, but assumes this constant to be zero (in Assumption A2).
Note that in certain works such as (Jiao et al., 2023), it is shown that the network size has to exponential in data dimension in order to achieve a small approximation error. However, in practice that would require an impractically large neural network size. In practice neural network size is of the same order as the data dimension. Thus for a fixed neural network size that we assume in this work, it makes sense to assume the approximation error as a constant.
@@
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function (Liu et al., 2022). Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models (Salimans & Ho, 2022; Liu et al., 2022). Note that in no prior works were such assumptions stated since they assumed access to the empirical risk minimizer.
@@
Usage of $p_{t_{0}}$ instead of $p_{0}$ in Theorem 1: We have shown that the estimated distribution $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in total variation (TV) to $p_{t_{0}}$, where $p_{t_{0}}$ denotes the data distribution pushed forward by $t_{0}$ steps of the forward process. We do not claim that $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in TV to the true data distribution $p_{0}$ (i.e., we do not bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$), because doing so would require additional assumptions on $p$. For example, Fu et al. (2024) (in Lemma D.5) assumes a sub-Gaussian data distribution to show that $\mathrm{TV}(p_{0},p_{t_{0}})\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)$. We also note that all other works listed in Table 1 similarly provide upper bounds on $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$, not on $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$.
@@
The bounds on $\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})$ and $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ follow from Lemma B.4 of Gupta et al. (2024) and Proposition 4 of Benton et al. (2024), respectively to get
$\mathrm{TV}((p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (16)
Note that we have used results from Gupta et al. (2024) and Benton et al. (2024) which assume a bounded second moment for the data distribution. This is satisfied by Assumption 1. Now from lemma 4, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ is upper bounded as follows
@@
In order to upper bound, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ we denote $A(k)$ as
@@
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Existing analyses of statistical errors, such as those given in Shalev-Shwartz & Ben-David (2014), only work when the loss function is bounded. This is not the case for diffusion models. Thus, we use a novel analysis that uses the conditional normality of the score function as well as the bounded second moment property of the data variable in Assumption 1 to obtain the upper bound on the statistical error. The details of the analysis are given in Appendix C.1.

### v6

@@ Abstract
Diffusion models have demonstrated state-of-the-art performance across vision, language, and scientific domains. Despite their empirical success, prior theoretical analyses of the sample complexity suffer from poor scaling with input data dimension or rely on unrealistic assumptions such as access to exact empirical risk minimizers. In this work, we provide a principled analysis of score estimation, establishing a sample complexity bound of $\mathcal{O}(\epsilon^{-4})$. Our approach leverages a structured decomposition of the score estimation error into statistical, approximation, and optimization errors, enabling us to eliminate the exponential dependence on neural network parameters that arises in prior analyses. It is the first such result that achieves sample complexity bounds without assuming access to the empirical risk minimizer of score function estimation loss.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks (ulhaq2022efficient; bansal2023universal), text generation (li2022diffusion), sequential data modeling (tashiro2021csdi), reinforcement learning and control (zhu2023diffusion), and life sciences (jing2022torsional; malusare2024improving). For a more comprehensive exposition of applications, we refer readers to survey paper (chen2024overview).
@@
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. Some of the key works studying the sample complexity are summarized in Table 1. A key limitation of sample complexity analyses of diffusion models done thus far is the lack of the presence of finite-time sample complexity results under reasonable assumptions. This makes the theoretical analysis of diffusion models fall short of other machine learning areas such as reinforcement Learning (kumar2023sample; gaur2024closinggapachievingglobal), bi-level-optimization (grazzi2023bilevel; gaur2025sample) and graphical models (8598839; tran2019sample). In this work we aim to bridge that gap and obtain a sample complexity results on the same footing as results from the aforementioned areas. The iteration complexity or convergence has been studied in li2024sharp; benton2024nearly; li2024adapting; huang2024denoisingdiffusionprobabilisticmodels; dou2024theory; liang2025low; liang2025broadening, while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function.
@@
We note that works such as zhang2024minimax; wibisono2024optimal; oko2023diffusion; chen2023score have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, guptaimproved improved upon this by obtaining $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity without exponential dependence on data dimension. [^1] We note that guptaimproved claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix A.. However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss, a significant restriction that was explicitly highlighted as an open problem in guptaimproved itself. While this assumption is present in all prior works, it is an unrealistic assumption regardless.
@@
zhang2024minimax $\tilde{O}\left(\epsilon^{-d}\right)$ Yes
wibisono2024optimal $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes
oko2023diffusion $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
chen2023score $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
guptaimproved1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes
@@
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. For further details on how the sample complexity bounds are derived for guptaimproved, see Appendix A.
@@
Unconditional and Conditional Diffusion Models. Diffusion models have emerged as leading frameworks across vision, audio, and scientific domains. Foundational works such as sohl2015deep and ho2020denoising introduced and refined Denoising Diffusion Probabilistic Models (DDPMs), enabling high-quality sample generation. Subsequent advances include improved noise schedules (nichol2021improved), score-based SDE formulations (song2021scorebased), and efficient latent-space generation via Latent Diffusion Models (LDMs) (rombach2022high). Conditional diffusion models extend these techniques for guided generation tasks, with applications in time-series tashiro2021csdi, speech huang2022fastdiff, and medical imaging dorjsembe2023conditional. Conditioning mechanisms range from classifier-based dhariwal2021diffusion to classifier-free guidance ho2022classifier, which enabled text-to-image models like Imagen saharia2022image and Stable Diffusion rombach2022high. Recent innovations focus on adaptive control castillo2025adaptive, compositionality liu2023more, and multi-modal conditioning avrahami2022blended.
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in chen2022sampling showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works li2024sharp; benton2024nearly; li2024adapting; huang2024denoisingdiffusionprobabilisticmodels; dou2024theory; liang2025low; liang2025broadening. In particular, benton2024nearly; li2024sharp establishes iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates pmlr-v235-li24ad; liang2025broadening; dou2024theory. Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in li2024adapting; huang2024denoisingdiffusionprobabilisticmodels; liang2025low. In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the errors introduced by the neural network approximation, data sampling, and optimization are also accounted for.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in chen2023score; zhang2024minimax; wibisono2024optimal; oko2023diffusion. However, these bounds are exponential in the data dimension. Further, authors of guptaimproved use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
@@ 2 Preliminaries and Problem Formulation
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process (oksendal2003stochastic), defined by the following SDE
@@
This is typically achieved using stochastic time-reversal theory (anderson1982reverse), which yields a corresponding reverse-time SDE as follows.
@@
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions. In practice, the backward process is run up to a fixed time point $t_{0}$ known as the early stopping time and not $t=0$. This is done in order to improve performance and training speed (lyu2022accelerating; favero2025bigger).
The continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework (ho2020denoising), where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix B. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
@@ 3 Sample Complexity of Diffusion Models
Some works that analyze the convergence of score-based diffusion models, such as chen2022sampling; oko2023diffusion, assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to have bounded second moment, making our results applicable to a significantly broader class of distributions.
@@
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses (liu2022loss). Prior works such as guptaimproved and block2020generative implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in guptaimproved and the definition of $\hat{f}$ in Theorem 13 of block2020generative). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.
@@
This error is independent of the sampling algorithm, and describes the error due to neural network parametrization. In learning theory, it is common to treat the approximation error of a model class as a constant so that analyzes can focus on the estimation/ optimization terms dependent on the sample. This convention appears in standard excess-risk decompositions for fixed hypothesis classes (shalev2014understanding). In PAC-Bayesian analyses, approximation errors are denoted by a constant once the class is fixed (mai2025pacbayesianriskboundsfully). In (NTK/RKHS) analyses of neural networks, where it is assumed the target function lies in, or is well approximated by the specified function class, the misspecification error is represented as a constant term (bing2025kernel). In reinforcement learning algorithm analysis such as policy gradient, a task-dependent “inherent Bellman” or function-approximation error that remains constant while deriving performance rates (mondal2024improved; fusingle; gaur2024closinggapachievingglobal; ganeshsharper). Note that guptaimproved also makes the same assumption implicitly, but assumes this constant to be zero (in Assumption A2).
Note that in certain works such as (jiao2023deep), it is shown that the network size has to exponential in data dimension in order to achieve a small approximation error. However, in practice that would require an impractically large neural network size. In practice neural network size is of the same order as the data dimension. Thus for a fixed neural network size that we assume in this work, it makes sense to assume the approximation error as a constant.
@@
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function (liu2022loss). Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models (salimans2022progressive; liu2022loss). Note that in no prior works were such assumptions stated since they assumed access to the empirical risk minimizer.
@@
Assume that the data distribution satisfies Assumption 1, the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2, 3,4 for all $k\in[0,K]$ and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$. Further assume
@@
Usage of $p_{t_{0}}$ instead of $p_{0}$ in Theorem 1: We have shown that the estimated distribution $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in total variation (TV) to $p_{t_{0}}$, where $p_{t_{0}}$ denotes the data distribution pushed forward by $t_{0}$ steps of the forward process. We do not claim that $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in TV to the true data distribution $p_{0}$ (i.e., we do not bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$), because doing so would require additional assumptions on $p$. For example, fu2024unveilconditionaldiffusionmodels (in Lemma D.5) assumes a sub-Gaussian data distribution to show that $\mathrm{TV}(p_{0},p_{t_{0}})\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)$. We also note that all other works listed in Table 1 similarly provide upper bounds on $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$, not on $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$.
@@
The bounds on $\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})$ and $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ follow from Lemma B.4 of guptaimproved and Proposition 4 of benton2024nearly, respectively to get
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (16)
Note that we have used results from guptaimproved and benton2024nearly which assume a bounded second moment for the data distribution. This is satisfied by Assumption 1. Now from lemma 4, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ is upper bounded as follows
@@
In order to upper bound $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$, we denote $A(k)$ as
@@
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Existing analyses of statistical errors, such as those given in shalev2014understanding, only work when the loss function is bounded. This is not the case for diffusion models. Thus, we use a novel analysis that uses the conditional normality of the score function as well as the bounded second moment property of the data variable in Assumption 1 to obtain the upper bound on the statistical error. The details of the analysis are given in Appendix C.1.